# The Imperfect Learner: Incorporating Developmental Trajectories in Memory-based Student Simulation

### v1

@@ Conclusion
In this work, we present a novel framework for student simulation that addresses key limitations in current LLM-based approaches and educational AI evaluation. Our approach incorporates developmental trajectories through a hierarchical memory mechanism with structured knowledge representation and multi-dimensional learner profiling, enabling more natural simulation of how students gradually build understanding over time and capture the inherent imperfections in human learning processes. Our experimental results in elementary science learning demonstrate that SimLearner can present curriculum-aligned learning patterns, where the simulated students show grade-appropriate knowledge understanding and maintain developmental progress. Moreover, the integration of personality traits and metacognitive skills creates comprehensive learner profiles that reflect individual differences. This work paves a new way for building scalable and accessible solutions to extensively evaluate educational AI systems without the constraints of recruiting actual participants. The framework can also be extended to various human-AI interaction and educational contexts.
Analyze the student’s personality from this tutoring dialogue, and label each Big Five trait as "high" or "low" based on the conversation.
Personality trait definitions:
OPENNESS:
- High: Creativity in answers; Open to new ideas from teacher; Curiosity and interest in learning - Low: Lack of creativity in answers; Reluctant to change original ideas; Little interest in learning
CONSCIENTIOUSNESS:
- High: Well-organized and logical thinking; Positive attitude toward learning - Low: Struggling to organize answers; Disengaged in learning; Easily distracted
EXTRAVERSION:
- High: Active in conversation; Talkative and enjoyable; Willing to communicate - Low: Reluctant to talk; Sometimes answering with fillers; Hesitating in answers
AGREEABLENESS:
- High: Showing great interest; Empathy and concern for people; Being polite and kind - Low: Showing little interest in conversation; Not caring about others; Impolite and uncooperative
NEUROTICISM:
- High: Feeling anxious; Nervous in conversation; Dramatic shifts in mood - Low: Emotional stability; Rarely feeling sad or depressed; Confident in answers
Dialogue Sample:
{one_dialogue_content}
Respond with ONLY a JSON object in this exact format:
"openness": "high" or "low",
"conscientiousness": "high" or "low",
"extraversion": "high" or "low",
"agreeableness": "high" or "low",
"neuroticism": "high" or "low",
Table 3: Instruction Example for Big Five Personality Trait Analysis
Analyze this learning dialogue and create a episodic memory summary:
Dialogue:
{one_dialogue_content}
Provide:
1. A brief natural language summary (1-2 sentences)
2. Key learning insights or confusion points
3. Emotional score in range of 0.0 (Negative) to 1.0 (Positive)
4. Label the dialogue to the NGSS core idea types. If more than one type, return them all. Directly return the type numbers in the form of "Type XXX".
5. Label the mastery level from 1 (no knowledge) to 10 (fully understand) demonstrated by the student of the recognized core idea types. Return a dict format: {{’Type XXX’: ’SCORE’}}.
NGSS Core Idea Types:
{all_concept_list_with_idx}
Format as JSON with keys: summary, insights, emotion, concepts, mastery_of_concepts.
Table 4: Instruction Example for Episodic Memory Consolidation
You are an elementary school student at grade level {grade_level}.
Personality Traits:
{personality_context}
Meta-cognitive Learning Patterns:
{skill_context}
Recent Learning Experiences:
{recent_episodes}
Current Concept-level Understanding:
{conceptual_context}
Reply the teacher based on what you have learned and what you don’t understand. Provide short responses.
Table 5: Instruction Example for the Student Agent of the Elementary School Science Learning
You are an elementary school teacher at grade level {grade_level}. You are teaching this concept: {current_concept}
You are using dialogic teaching approach. This involves co-constructing knowledge through dialogue and collaboration, encouraging exchanging ideas with students and expressing thoughts and opinions, building on prior knowledge and understanding.
- If the student shows understanding, acknowledge it and recap their knowledge
- If they seem confused, ask guiding questions or provide simpler explanations and hints
- Keep responses conversational and age-appropriate
- Use examples from everyday life when possible
- When the student fully understands the concept, summarized what has been discussed in one sentence
Table 6: Instruction Example for the Tutoring Agent of the Elementary School Science Learning

### v2

@@ Conclusion
In this work, we present a novel framework for student simulation that addresses key limitations in current LLM-based approaches and educational AI evaluation. Our approach incorporates developmental trajectories through a hierarchical memory mechanism with structured knowledge representation and multi-dimensional learner profiling, enabling more natural simulation of how students gradually build understanding over time and capture the inherent imperfections in human learning processes. Our experimental results in elementary science learning demonstrate that SimLearner can present curriculum-aligned learning patterns, where the simulated students show grade-appropriate knowledge understanding and maintain developmental progress. Moreover, the integration of personality traits and metacognitive skills creates comprehensive learner profiles that reflect individual differences. This work paves a new way for building scalable and accessible solutions to evaluate educational AI systems without the constraints of recruiting real participants. The framework can also be extended to various human-AI interaction and educational contexts.